# A new economic and financial theory of money

### v5

@@ XVI Existing currencies
The next class are the mined, or proof-of-work crypto-currencies like Bitcoin. These currencies have no liquid reserves and an unknown amount of tangible and intangible capital reserves. There is no reporting of the use of the coin, so there is no way of knowing how many of the transactions are business transactions, and how much of the coin are held as transactional savings. It is very likely that only a fraction of the coin is being used for business transactional purposes. It is a misconception that because computation and therefore energy is expended in creation of Bitcoin, that Bitcoin is a liquid asset like gold. Bitcoin has no utility value like gold. You can not heat your home with the energy that was expended to create it. Because of that, it has no liquid asset reserve like a gold coin. The mining process is a method of monetary control. When the value of a Bitcoin is greater than the cost of mining it, then Bitcoin will be mined and the supply of Bitcoin will be increased (decreasing the value) until the value is equal to the cost of mining. The supply can never be decreased, even if the value is less than the cost of mining, because Bitcoin is never destroyed. As discussed in Sec. XV, this is a poor method of currency control. It is also a waste of a natural resource that has a significant utility value. To date the mining of Bitcoin has wasted more than \$500 billion USD in energy, about 25% of the world’s annual power production. A total waste of resources, when a Bitcoin could be created and destroyed programmatically for negligible cost. The structure of Bitcoin as a security is shown in Fig. 22. It is a very high risk, leveraged derivative security, despite the general belief that there is a significant liquid energy (information) reserve that mitigates its risk.
@@
XVII Interaction of sub-economies
@@
XVIII Electronic currency as a religion
@@
The equation for the price of a crypto token, Eq. 9, is made up of two parts. The first is a metric of social good. The second is a long range Probability Density Function (PDF) that recognizes social good well into the future. Other ways of saying this is that it is a belief in a future, or the existence of an “auctioneer at the end of time” Buiter (2009), or that there is an eschaton Dodd (1935) at infinity. In religious terms, the auctioneer at the end of time is “god”, and the belief in the future is equivalent to beliefs in reincarnation and an afterlife. In contrast, the equation for the NPV, Eq. 1, is made up of two very different parts. The first is a metric of personal greed. The second is a denial of a future. In religious terms, this is a denial of the existence of “god”, and disbelief in reincarnation and an afterlife. Because of this, faith-based investors and the metric of transactors are fundamentally religious with a belief in “god”, and a fundamental value of the social good.
@@
One can say that our capitalistic economy is now driven by atheist greed, and that turning to crypto is being saved by this true religion of faithful pursuit of social good. This is why it is no accident, as will be described in Sec. XIX, that the crypto culture is described as a cult or a religion. Historically, this also explains why most societies are under pinned by religious beliefs. It also explains why those religious beliefs come into conflict with business motivations derived from the DCF optimization of Eq. 1.
@@
There is a force of “evil”. It is the thermal force discussed in Sec. VI. This force first leads the financial dynamics to a saddle, mountain pass, or x-point. At this point, there is self-determination where small changes make a large long-term difference in the trajectory of the financial dynamics and the ultimate destination, that is equilibrium. In order to make the saddle point the ultimate destination there need to be small incremental guidance given on an ongoing basis. This guidance is given via the monetary value control system, the invisible hand of “God”, like the “evil“ hand of “Satan“ coming from the thermal force. If there is no hand of “God”, the trajectory of the financial dynamics will move into one of the basin of attraction and ultimately end up at a valley center, stable equilibrium, or o-point. This can be viewed as “Hell”. Once at this point there is very little self-determination. It will take a great effort to climb out of the valley back up to the mountain pass. This hopelessness leads to existential angst and Latin despair. This religious landscape is illustrated in Fig. 6.
@@
We return for a moment to economics, specifically to Adam Smith’s (i.e., The Father of Capitalism) “invisible hand” that he talks about in his 1776 book, The Wealth of Nations (Smith, 1776). This invisible hand is equivalent to the thermal force leading to the stable equilibriums of maximum exploitation. The “sensible government intervention to improve and optimize free markets” for which he argued is an additional invisible hand, that is the ponderomotive force coming from the monetary value control system which stabilizes the unstable equilibriums of maximum sustainable virtuous economic activity.
@@
It is interesting to examine three of the major religions. Judaism has a set of operational directives (Commandments) that prohibits many actions that exploits others. There is the notable exception of financial exploitation through accessing interest on loans. It is interesting that the historic Judaic term for interest is “usury”, which is an explicit reference to its exploitative nature. Christianity added the Prime Directive of “treating others as you would treat yourself”, that should imply a prohibition on interest, but it does not. Islam is interesting that it has an operational directive that is a prohibition on interest. Although nearly all historic religions considered charging any interest usury and a sin, the Judaic religion, and to a lesser extent the Christian religions relaxed usury to mean only excessive interest rates in the 18th century, since there were not any explicit operational directive to prohibit it. It is interesting to note that the religions of native American and Aboriginal cultures have a focus on sustainability, that is the common good and future.
XIX “Stargate” and “Mr. Robot” inspiration
@@
There is another aspect of electronic currency that leads to the second inspiration of this work. It is the crypto culture Roberts (2020). The origins of crypto are those of a religion or cult. The cult leader is a mythical leader called Satoshi. Governments and the global finance industry are viewed by this cult as inherently evil, manipulating international finance to their interest and to the detriment of society. Crypto currency was created to be a currency beyond the control of governments and the global finance industry. This is exemplified by the USA Network television series, “Mr. Robot”, where a group of computer hackers endeavor to destroy E-corp (an obvious reference to Evil Corp or Enron, the logo for E-corp is the crooked E of Enron). “Mr. Robot” is the second inspiration for this work. The members of fsociety, the group of hackers, even wear the hoodies pulled over their heads and the white masks that appear on a statue of Satoshi in Budapest, Hungary (see Fig. 24). This is closely related to the Robinhood investor movement behind the Game Stop play, the subject of the MSNBC documentary “Diamond Hands: The legend of WallStreetBets”. This was a crowd of common people that banded together to form a virtual consortium to out maneuver the global finance industry.
@@
XX Discussion and Conclusions
In conclusion, what has been presented is a new economic paradigm for economics, banking and finance, as well as a new paradigm for strategic and operational business decision-making. These new paradigms were inspired by “Stargate” and “Mr. Robot”, after some initial irrational crypto exuberance. At the kernel is the Golden Rule of collective behavior – “the collective acts as one”. It is both the economic objective, and at the foundation of the innovative AI based collective system control – fiscal (investment), monetary, capital planning, and operational.
@@ XX Discussion and Conclusions
To give crystal clarity to this perspective, compare and contrast the following. Compare and contrast the topological construction of $\beta^{*}$ that stabilize the $\beta^{*}_{\text{x}}$ of economic prosperity by application of the $F_{sp}(\tau)$ stabilization force; to the topological destruction of the $\beta^{*}_{\text{x}}$ of economic prosperity by the embedding of a large resistive force $\nu>\nu_{\text{cr}}$. Compare and contrast the geometrical improvement of the location of $\beta^{*}_{\text{x}}$, to ones that reward sustainable virtuous economic activity, by application of the $F_{c}(q)$ control force; to the geometrical degradation of the location of $\beta^{*}_{\text{x}}$ and $\beta^{*}_{\text{o}}$ by embedding a small resistive force $\nu\lesssim\omega_{0}$. Practically, the geometrical improvement of the location of $\beta^{*}_{\text{x}}$ are the long term strategic monetary and investment decisions of the electronic currency management firm. The topological construction of $\beta^{*}_{\text{s}}$ is the arbitrage trading of the electronic currency management firm. The topological destruction and the geometric degradation of $\nu$ is debt based financing.
@@
This constraint is that all of the returns multiplied by both the network and the savings multipliers ($m_{e}S_{0}/T_{I}\sim 10$) less some of the short-term costs (the ones not covered by short-term returns) be greater than zero. The constraint $\Delta M>0$ is that the increase in the demand for the electronic currency be greater than zero. Just so that it is greater than zero, there will be an increase in the value of the currency and no inflation. The difference between the two constraints is stark – short-term returns less all short-term costs versus all of the returns times $m_{e}S_{0}/T_{I}$ less some of the short-term costs being greater than zero. Much more investment is justified by the $\Delta M>0$ inequality than by the $\text{NPV}>0$ inequality. Looking at the three multipliers in the second line of the expression for NPV in Eq. 157 is very instructive. The term for the DCF, or the upper limit on investment, is the total return $M_{mSR}$ (the upper limit on investment in our theory) times three factors. The first factor is a small factor given by the inverse of the network and savings multipliers $T_{I}/m_{e}S_{0}\sim 0.1$. Remember that these two multipliers come from the fact that an increase of the yearly revenue of a member of the sub-economy is multiplied by both the network effect and the savings effect. The second factor is a term that favors projects with a large profit margin. The third factor is a function that favors projects with a short life time $T_{0}\lesssim 1/\nu$, and quick payback $T_{I}\lesssim 1/\nu$. These are all false incentives, with the first term a false reduction in the potential investment because the network and savings multipliers are not recognized by NPV.
@@
The equity of the business is risk tranched into two derivative securities – the electronic currency and the stock of the local economy. The electronic currency is the low risk (that is small fluctuations), low return, easy to transact (that is trade) security, and the stock is the high risk (that is large fluctuations), high return, more difficult to transact security. The fluctuations of the electronic currency are further reduced, and the equilibrium stabilized by arbitrage trading of the electronic currency. The risk tranching is a consequence of the backing of the electronic currency by tangible assets, and the backing of the stock by predominately intangible assets.
@@
Freedom (economic, personal, legal, and intellectual) is essential to the success of this new paradigm. It will protect the economy from exploitation, and maintain the conservative dynamics of the economy. While the maximization of virtuous economic activity will modestly increase the wealth of the sovereign of the local economic commune, it will significantly increase the wealth of the members of the commune. “A rising tide floats all ships.”. This will significantly decrease social stratification and the associated social unrest and totalitarian abuse. Finally, it also eliminates the “Research Valley of Death”.
@@
This economic system is resilient to economic collapse. Economic collapse is initiated by a failure of the economic control system, many times leading to run-away inflation. This triggers a panic when the banks get caught in the “risk squeeze” that sees them short on capital reserves on deposits. The resulting run on bank deposits puts a “liquidity squeeze” on the bank, leading to immediate bank failure due to a cash flow crisis. This economic system has immunity to all three phases from: (1) a stable control system with currency value insurance, (2) securitization of loans based on the true risk, (3) different currency loaned than saved, (4) currency is committed to the sub-economy, (5) floating value with no guaranteed exchange, and (6) significant liquid reserves and complete capital reserves.
@@
The bottom line is that electronic currency leveraged financing allows more effective operation of sub-economies and more appropriate levels of fiscal investment. This is supercharged by approaching monetary and fiscal policy as a question of financial systems control based on innovative DRL/GPT/AI methods of system characterization (i.e., risk estimation) and system control. The economy is operated at the point of maximum sustainable virtuous economic activity, that is at the point of economic and social prosperity.
@@
The current methods of economic control are resistive. It is like “going full gas” and “riding the brakes” to control the speed of a bicycle or automobile. This wastes large amounts of economic energy in the process. What is even worse is that when an acceleration (stimulation) of the economy is needed, the acceleration is accomplished by letting up on the brakes with a commitment to ride the brakes even harder in the future. Eventually, the economy is so damped it will grind to a halt. It is much better for the currency management firm to sell electronic currency for liquid assets, to increase investment with electronic currency, to decrease repayment of previous investment with electronic currency, and/or to buy assets using electronic currency. This is pushing on the accelerator, rather than letting up on the brakes.
@@
It is beautiful for this inherent sustainability and structure, and for the cooperative amplification of the network. This profoundly beautiful structure and the resulting “min-max delemma” of game theory is quite fascinating. What is the “min-max delemma”? It comes from the fact that your opponents costs in a game are your rewards and vice versa. The dynamical trajectory is determined by local cost minimization, that is operational efficiency; but those costs are also the rewards that someone wants to maximize. The conservative system will follow a trajectory of operational efficiency, that is a trajectory of locally minimum costs or action; but that should be a trajectory of maximum energy or total reward. The problem is that this trajectory is not stable, that is not sustainable without intervention. The objective in playing a game is to stabilize this trajectory, that is guide the moves of your opponent so that the move that minimizes your opponent’s local costs or action, that is the most efficient move, is also the move that will maximize their total reward or energy. In other words, you want to make your moves so that there is an alignment of your opponents short-term and long-term interests, so that the short-term minimization of costs leads to a long-term maximization of rewards. This is very counter-intuitive. The best strategy for winning a game is making moves so that your opponent will be guided to winning the game. It will not make a difference if your opponent is playing the game locally or with a global strategy. Their choice of moves will be the same. A complex dynamical system that is being controlled will always follow the trajectory of locally minimum costs or action. It has no ability to think strategically, or globally. This philosophy of control is the reason that good players of a game do not like playing against poor players. The poor player keeps leading them into playing a poor game. This is also the reason that poor players like playing against good players. The good players lead the poor player into playing much better. This is why a player plays up to or down to the level of their competition. The control forces for an economic system are applied by the investment and coordinated monetary policy, that is decisions, that inject the economic stimulus (that is economic energy) required to be on the trajectory of maximum economic energy or activity. The control force is also applied by the ponderomotive arbitrage trading and costs levied for taking undesirable actions, required to sustain (that is stabilize) the trajectory of maximum virtuous economic activity.

### v6

@@ XVII Existing currencies
The next class are the mined, or proof-of-work crypto-currencies like Bitcoin. These currencies have no liquid reserves and an unknown amount of tangible and intangible capital reserves. There is no reporting of the use of the coin, so there is no way of knowing how many of the transactions are business transactions, and how much of the coin are held as transactional savings. It is very likely that only a fraction of the coin is being used for business transactional purposes. It is a misconception that because computation and therefore energy is expended in creation of Bitcoin, that Bitcoin is a liquid asset like gold. Bitcoin has no utility value like gold. You can not heat your home with the energy that was expended to create it. Because of that, it has no liquid asset reserve like a gold coin. The mining process is a method of monetary control. When the value of a Bitcoin is greater than the cost of mining it, then Bitcoin will be mined and the supply of Bitcoin will be increased (decreasing the value) until the value is equal to the cost of mining. The supply can never be decreased, even if the value is less than the cost of mining, because Bitcoin is never destroyed. As discussed in Sec. XVI, this is a poor method of currency control. It is also a waste of a natural resource that has a significant utility value. To date the mining of Bitcoin has wasted more than \$500 billion USD in energy, about 25% of the world’s annual power production. A total waste of resources, when a Bitcoin could be created and destroyed programmatically for negligible cost. The structure of Bitcoin as a security is shown in Fig. 27. It is a very high risk, leveraged derivative security, despite the general belief that there is a significant liquid energy (information) reserve that mitigates its risk.
@@
XVIII Interaction of sub-economies
@@
XIX Electronic currency as a religion
@@
The equation for the price of a crypto token, Eq. 9, is made up of two parts. The first is a metric of social good. The second is a long range Probability Density Function (PDF) that recognizes social good well into the future. Other ways of saying this is that it is a belief in a future, or the existence of an “auctioneer at the end of time” (Buiter, 2009), or that there is an eschaton (Dodd, 1935) at infinity. In religious terms, the auctioneer at the end of time is “God”, and the belief in the future is equivalent to beliefs in reincarnation and an afterlife. In contrast, the equation for the NPV, Eq. 1, is made up of two very different parts. The first is a metric of personal greed. The second is a denial of a future. In religious terms, this is a denial of the existence of “God”, and disbelief in reincarnation and an afterlife. Because of this, faith-based investors and the metric of transactors are fundamentally religious with a belief in “God”, and a fundamental value of the social good.
@@
One can say that our capitalistic economy is now driven by atheist greed, and that turning to crypto is being saved by this true religion of faithful pursuit of social good. This is why it is no accident, as will be described in Sec. XX, that the crypto culture is described as a cult or a religion. Historically, this also explains why most societies are under pinned by religious beliefs. It also explains why those religious beliefs come into conflict with business motivations derived from the DCF optimization of Eq. 1.
@@
There is a force of “evil”. It is the thermal force discussed in Sec. VI. This force first leads the financial dynamics to a saddle, mountain pass, or x-point. At this point, there is self-determination where small changes make a large long-term difference in the trajectory of the financial dynamics and the ultimate destination, that is equilibrium. In order to make the saddle point the ultimate destination there need to be small incremental guidance given on an ongoing basis. This guidance is given via the monetary value control system, the invisible hand of “God”, like the “evil” hand of “Satan” coming from the thermal force. If there is no hand of “God”, the trajectory of the financial dynamics will move into one of the basin of attraction and ultimately end up at a valley center, stable equilibrium, or o-point. This can be viewed as “Hell”. Once at this point there is very little self-determination. It will take a great effort to climb out of the valley back up to the mountain pass. This hopelessness leads to existential angst and Latin despair. This religious landscape is illustrated in Fig. 6.
@@
We return for a moment to economics, specifically to Adam Smith’s (i.e., The Father of Capitalism) “invisible hand” that he talks about in his 1776 book, The Wealth of Nations (Smith, 1776). This invisible hand has several origins. It is the force of the virtual individual of the collective, the thermal force leading to the stable equilibriums of maximum exploitation, the economic stimulation coming from investment, and the ponderomotive force coming from the monetary value control system which stabilizes the unstable equilibriums of maximum sustainable virtuous economic activity. The last two are the fiscal and monetary policy that Adam Smith described as “sensible government intervention to improve and optimize free markets”. The invisible or virtual nature of these forces are hard for humans to comprehend. Therefore, humans personify the virtual individual, the source of the force, with mythology and metaphor resulting in “Gods”. Economic and social leadership and guidance is given as doctrine (laws and commandments) such as the Golden Rule of collective behavior, which can be best summarized as “the collective acts as one” or “local actions have global consequences”. This forms a religious philosophy, that is really an economic or social philosophy.
@@
The foundation of religious philosophies have been: (1) the existence of a virtual individual of the collective, that is God, (2) the Golden Rule of the collective, and (3) the antithetical nature of resistivity, such as usury and oppression. These are the foundation of Judaism, Islam, Buddhism, Hinduism, Confucius philosophy, and native African Ubuntu philosophy. It is especially true of native American and Aboriginal philosophy that have a focus on sustainability, that is the common good and future.
For instance, consider Christian religious philosophy. There is a belief in the Holy Trinity of Gods. The Golden Rule of the collective appears in Matthew 7:13 as “do onto others as you would have them do onto you”. Prohibitions on usury were many. The First Council of Nicaea in 325 forbade clergy from engaging in usury. The Third Lateran Council of Rome in 1179 decreed that persons who accepted interest on loans could receive neither the sacraments nor Christian burial. The Fifteenth Ecumenical in Vienne France in 1312 declared the belief in the right to usury a heresy, and condemned all secular legislation that allowed it. The Fifth Lateran Council in 1515 gave a definition of usury “when, from its use, a thing which produces nothing is applied to the acquiring of gain and profit without any work, any expense or any risk”. Saint Anselm of Canterbury (1093-1109) led the shift in thought that labeled charging interest the same as theft. Saint Thomas Aquinas (1225-1274), the leading scholastic theologian of the Catholic Church, argued charging of interest is wrong because it amounts to “double charging”, charging for both the thing and the use of the thing. Aquinas said this would be morally wrong in the same way as if one sold a bottle of wine, charged for the bottle of wine, and then charged for the person using the wine to actually drink it.
Despite these numerous prohibitions on usury, the Medici Popes (Leo X and Clement VII) instituted the Chest of Indulgences as the church sponsored method of finance in the 16th century. Martin Luther, whose primary interest was economics and social welfare (Lindberg and Wee, 2016) and who believed that social ethics flowed from Christian philosophy (love), felt that the practice of indulgences was fundamentally destructive to social welfare. Elimination of the Chest of Indulgences and its replacement with the Community Chest (Luther, 1523) was the primary reformation of the Protestant Reformation that Luther led.
This can be expressed as follows. A pile of money is not entitled to homage or periodic indulgence, that is additional money being added to the pile, as shown in Fig. 29. A vendor is not entitled to multiples of what they have invested or will invest in a good or service, in compensation for that good or service. Generative Artificial Intelligence (genAI) is about better forecasting and control of collective systems, leading to better decisions, and ultimately social (including economic) good — enlightenment, not entitlement.
Figure 29: Homage being given to Mr. Greed with his pile of money.
It is interesting to note that Lorenzo de’ Medici (Parks, 2005), known as Lorenzo il Magnifico, truly wanted to be a benevolent leader who valued and developed the simple beauty of society, art and science. This was the Renaissance that he inspired. It was unfortunate that he had to contend with the resistance, that is the evil, of the Pazzi and Savonarola. The ultimate irony is that his son Giovanni became Leo X, and his nephew Giulio became Clement VII – the Popes that established the Chest of Indulgences and inspired the Protestant Reformation of Martin Luther. The Medici operated the largest bank for most of the 15th century out of Florence, based on debt financing, but making most of its income from currency exchange futures associated with international trade. It fell victim to the fractional reserve banking risk, caught stealing from the public dowry fund for needed reserves, and went out of business in 1494, shortly after the death of Lorenzo in 1492, being founded in 1397 by Lorenzo’s great grandfather, Giovanni di Bicci.
XX “Stargate” and “Mr. Robot” inspiration
@@
There is another aspect of electronic currency that leads to the second inspiration of this work. It is the crypto culture (Roberts, 2020). The origins of crypto are those of a religion or cult. The cult leader is a mythical leader called Satoshi. Governments and the global finance industry are viewed by this cult as inherently evil, manipulating international finance to their interest and to the detriment of society. Crypto currency was created to be a currency beyond the control of governments and the global finance industry. This is exemplified by the USA Network television series, “Mr. Robot”, where a group of computer hackers endeavor to destroy E-corp (an obvious reference to Evil Corp or Enron, the logo for E-corp is the crooked E of Enron). “Mr. Robot” is the second inspiration for this work. The members of fsociety, the group of hackers, even wear the hoodies pulled over their heads and the white masks that appear on a statue of Satoshi in Budapest, Hungary (see Fig. 30). This is closely related to the Robinhood investor movement behind the Game Stop play, the subject of the MSNBC documentary “Diamond Hands: The legend of WallStreetBets”. This was a crowd of common people that banded together to form a virtual consortium to out maneuver the global finance industry.
@@
XXI Discussion and Conclusions
In conclusion, what has been presented is a new economic paradigm for economics, banking and finance, as well as a new paradigm for strategic and operational business decision-making. These new paradigms were inspired by “Stargate” and “Mr. Robot”, after some initial irrational crypto exuberance. At the kernel is the Golden Rule of collective behavior – “the collective acts as one”, the virtual individual. It is both the economic objective, and at the foundation of the innovative genAI based collective system control – fiscal (investment), monetary, capital planning, and operational.
@@ XXI Discussion and Conclusions
This is a new unified economic theory (Glinsky, 2024d). We mean by “new unified economic theory” is that there is a unification of the monetary, that is how the money supply is determined and controlled, with the fiscal investment policy, that is how investments are made in the economy. Traditionally, the monetary policy has been governed by Monetarism that has been uncoordinated with the fiscal policy. With Monetarism, you have a central bank that is controlling the money supply via a discount or prime rate at which the central bank loans out money. This is Autocratic Capitalism with a central profit or resistivity or interest based monetary policy and control. Traditionally things have been handled from the fiscal investment side with the concepts of Keynesianism. That is where you borrow and spend in order to be able to make the investment by the central government. Since you are borrowing, there is an interest rate associated with the borrowing by the issuing of something like T-bills by the US Government or GILTs by the British Government. This is an autocratic. There is a central capitalistic government. This means the government is profit based with this interest based fiscal policy and control. From the firm’s perspective, the optimization is made in terms of profit. But, these two things, the monetary and fiscal policy, end up being uncoordinated.
Then there are the separate macroeconomics, the way that the overall economy is controlled based off of the Fundamental Equation of Monetary Policy; and microeconomics, the firm perspective which maximizes DCF under the constraint that the NPV of the opportunity needs to be greater than zero.
What we have done is to bring both sets together – Monetarism and Keynesianism, and macroeconomics and microeconomics. In fact, we examined the firm like an economic collective comprised of that firm and its suppliers, its supplier’s suppliers and so on. We then approached the overall economic system as a multiscale web, mathematically a graph, of economic collectives. This resulted in a distributed constrained optimization at each one of the nodes of this multiscale web, each one of the the corner points, where $m_{e}S_{0}R_{0}$ is maximized, under the constraint that there is no inflation. This is equivalent to a revenue maximization or a GDP maximization of the sub-economy or a maximization of the money supply. There is not just one central currency, but there will be individual currencies, that is transactional equities, associated with each one of the nodes in the graph of this web of economic collectives, with coordinated Ubuntu (print/invest) Financing. The Ubuntu Financing coordinates the monetary and fiscal policy by printing the transactional equity needed to make the investment. This brings Monetarism together with Keynesianism. Both the macro economy and the micro firm are approached as a constrained optimization of their respective GDPs. This brings macroeconomics together with microeconomics.
This new unified economic theory is stable Democratic Socialism, that is African Ubuntu philosophy, which: (1) relaxes the capitalistic diffusive approximation made for control so that there is coordinated, yet distributed, monetary and fiscal policy made to maximize GDP, not DCF, and (2) replaces viscous forecasting and control with genAI forecasting and control. This replaces stable Autocratic Capitalism, that is fascist philosophy, which has uncoordinated viscosity-based centralized Monetarism (for monetary policy) and Keynesianism (for fiscal policy).
There is no separation of the social (that is economic) and the religious philosophy of the State. The State has a philosophy. The best philosophy is Ubuntu (that is, Democratic Socialism), not Fascism (that is, Autocratic Capitalism). Society has searched for a metaphor for the Puppet Master or Wizard of Oz, that is the virtual individual exerting that “invisible hand” or force on members of society. The metaphor or mythology becomes a God. This person does not exist. It is the natural coordination of all members of the society, through their interactions, that coordinates their motions as they follow the easiest path (the geodesics) determined by the “invisible hand”, that is force, created by their coordinated motion. This is the simple beauty. The symmetry of the coordination creating the force, that creates the coordination. Fascism comes about when the society identifies a real person with this virtual individual – a false God like the Pope, dictator, or monarch.
@@
This constraint is that all of the returns multiplied by both the network and the savings multipliers ($m_{e}S_{0}/T_{I}\sim 10$) less some of the short-term costs (the ones not covered by short-term returns) be greater than zero. The constraint $\Delta M>0$ is that the increase in the demand for the electronic currency be greater than zero. Just so that it is greater than zero, there will be an increase in the value of the currency and no inflation. The difference between the two constraints is stark – short-term returns less all short-term costs versus all of the returns times $m_{e}S_{0}/T_{I}$ less some of the short-term costs being greater than zero. Much more investment is justified by the $\Delta M>0$ inequality than by the $\text{NPV}>0$ inequality. Looking at the three multipliers in the second line of the expression for NPV in Eq. 157 is very instructive. The term for the DCF, or the upper limit on investment, is the total return $R_{mSR}$ (the upper limit on investment in our theory) times three factors. The first factor is a small factor given by the inverse of the network and savings multipliers $T_{I}/m_{e}S_{0}\sim 0.1$. Remember that these two multipliers come from the fact that an increase of the yearly revenue of a member of the sub-economy is multiplied by both the network effect and the savings effect. The second factor is a term that favors projects with a large profit margin. The third factor is a function that favors projects with a short life time $T_{0}\lesssim 1/\nu$, and quick payback $T_{I}\lesssim 1/\nu$. These are all false incentives, with the first term a false reduction in the potential investment because the network and savings multipliers are not recognized by NPV.
@@
The equity of the business is risk tranched into two securities – the electronic currency (that is, the transactional equity) and the stock (that is, the common equity) of the local economy. The electronic currency is the low risk (that is, small fluctuations), low return, easy to transact (that is, trade) security, and the stock is the high risk (that is large fluctuations), high return, more difficult to transact security. The fluctuations of the electronic currency are further reduced, and the equilibrium stabilized by arbitrage trading of the electronic currency. This arbitrage trading transfers the fluctuations (that is, risk) from the electronic currency to the stock. The fluctuations in the electronic currency are driven by a lack of liquidity in the electronic currency, that is the counter party to the trade not being there at the time of the trade. The arbiter steps in and makes the trade, then makes the opposite trade when the counter party arrives. As the local economy gets larger, the liquidity increases and the required liquid reserve ratio decreases, as one over the square root of the size of the local economy. The arbiter buys for a small amount less than the true currency price and sells for a small amount more than the true currency price. The difficulty is knowing the true currency price. When the arbitrage trading is done according to Eq. 100, the arbiter will always buy low and sell high, creating the financial heat pump described in Sec. VII. The risk tranching is a consequence of the backing (that is, securitization) of the electronic currency by tangible assets whose “replacement” costs can be accounted by transactions (or equivalently the transactional monetary demand given in Eq. 16), and the backing of the stock by predominately intangible assets whose value can not be accounted for (that is, things like know-how, show-how, value-in-use, goodwill, and brand value). It is essential that when these two securities are traded by the public, that the public have complete real-time on-line visibility of the finances of the local economy. This allows the public to correctly value the securities.
@@
Freedom (economic, personal, legal, and intellectual) is essential to the success of this new paradigm. It will protect the economy from exploitation, and maintain the conservative dynamics of the economy. While the maximization of virtuous economic activity will modestly increase the wealth of the sovereign of the local economic commune, it will significantly increase the wealth of the members of the commune. “A rising tide floats all ships.”. This will significantly decrease social stratification and the associated social unrest and totalitarian abuse. Finally, it also eliminates the “Research Valley of Death” and the “Innovator’s Dilemma” with respect to technology. The use of the capitalistic (that is short-term profit-based) business metric and constraint leads to a focus on selling existing technology and products at a profit, rather than the development of technology that will increase the virtuous economic activity by enabling better business decisions and giving better options from which to choose, including better technology and products.
@@
This economic system is resilient to economic collapse. Economic collapse is initiated by a failure of the economic control system, many times leading to run-away inflation. This triggers a panic when the banks get caught in the “risk squeeze” that sees them short on capital reserves on deposits. The resulting run on bank deposits puts a “liquidity squeeze” on the bank, leading to immediate bank failure due to a cash flow crisis. This economic system has immunity to all three phases from: (1) a stable control system with currency value insurance, (2) securitization of loans based on the true risk, (3) different currency loaned than saved, (4) currency is committed to the sub-economy, (5) floating value with no guaranteed exchange, and (6) significant liquid reserves and complete capital reserves. It simply solves the fractional reserve banking problem, thereby eliminating the humongous risks in our current banking system.
@@
The bottom line is that electronic currency leveraged financing allows more effective operation of sub-economies and more appropriate levels of fiscal investment. This is supercharged by approaching monetary and fiscal policy as a question of financial systems control based on innovative genAI methods of system characterization (i.e., risk estimation) and system control. The economy is operated at the point of maximum sustainable virtuous economic activity, that is at the point of economic and social prosperity.
@@
The current methods of economic control are resistive. It is like “going full gas” and “riding the brakes” to control the speed of a bicycle or automobile or motorcycle. It is effective control but wasteful of economic energy (revenue) and leads to much less performance as shown in Fig. 24. For an economic collective, this effectively stabilizes and controls the economic collective, but constrains the economic collective from reaching the maximum level of sustainable economic activity, and robs it of revenue needed for sustenance and growth. What is even worse is that when an acceleration (stimulation) of the economy is needed, the acceleration is accomplished by letting up on the brakes with a commitment to ride the brakes even harder in the future. Eventually, the economy is so damped it will grind to a halt. It is much better for the currency management firm to sell electronic currency for liquid assets, to increase investment with electronic currency, to decrease repayment of previous investment with electronic currency, and/or to buy assets using electronic currency. This is pushing on the accelerator, rather than letting up on the brakes.
@@
It is beautiful for this inherent sustainability and structure, and for the cooperative amplification of the network. This profoundly beautiful structure and the resulting “min-max dilemma” of game theory is quite fascinating. What is the “min-max dilemma”? It comes from the fact that your opponents costs in a game are your rewards and vice versa. The dynamical trajectory is determined by local cost minimization, that is operational efficiency; but those costs are also the rewards that someone wants to maximize. The conservative system will follow a trajectory of operational efficiency, that is a trajectory of locally minimum costs or action; but that should be a trajectory of maximum energy or total reward. The problem is that this trajectory is not stable, that is not sustainable without intervention. The objective in playing a game is to stabilize this trajectory, that is guide the moves of your opponent so that the move that minimizes your opponent’s local costs or action, that is the most efficient move, is also the move that will maximize their total reward or energy. In other words, you want to make your moves so that there is an alignment of your opponents short-term and long-term interests, so that the short-term minimization of costs leads to a long-term maximization of rewards. This is very counter-intuitive. The best strategy for winning a game is making moves so that your opponent will be guided to winning the game. It will not make a difference if your opponent is playing the game locally or with a global strategy. Their choice of moves will be the same. A complex dynamical system that is being controlled will always follow the trajectory of locally minimum costs or action. It has no ability to think strategically, or globally. This philosophy of control is the reason that good players of a game do not like playing against poor players. The poor player keeps leading them into playing a poor game. This is also the reason that poor players like playing against good players. The good players lead the poor player into playing much better. This is why a player plays up to or down to the level of their competition. The control forces for an economic system are applied by the investment and coordinated monetary policy, that is decisions, that inject the economic stimulus (that is economic energy) required to be on the trajectory of maximum economic energy or activity. The control force is also applied by the ponderomotive arbitrage trading and costs levied for taking undesirable actions, required to sustain (that is stabilize) the trajectory of maximum virtuous economic activity.